Imagine that you do 10 pull-ups in 10 seconds. What is the average power supplied by your arm muscles when moving your body?



1
Expert's answer
2021-03-16T11:20:03-0400


The work done against the force of gravity (when going up or down) is "mg\Delta y". Therefore, the work done for a complete repetition is "2mg\Delta y". Let's suppose that the mass of the person is 80 kg and the distance "\Delta y" equals 60 cm. Also, let suppose that the arms comprise 10% of the body mass and are not included in the moving mass. Then, we can calculate the average power supplied by the arm muscles when moving the person's body:


"P_{avg}=\dfrac{W}{t},""P_{avg}=\dfrac{10\cdot2\cdot(1-0.1)\cdot80\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot0.6\ m}{10\ s}=847\ W."